A boat goes 24 km upstream and 28 km downstream in 6 hours . It goes 30 km upstream and 21 km downstream in 6 hours 30 minutes . Find the speed of the boat in still water and also the speed of the stream
Answer by Paul(988)

Differnce between upstream:
6km
Difference between downstream:
7km.
difference between time:
0.15minutes:
Let the speed of the boat be x
Let the speed of the stream be y:
0.15(x-y)=6
0.15x-0.15y=6
0.15y=-6+0.15x
y=-40+x (subsituion)

0.15(x+y)=7
0.15x+0.15y=7
Subsitute for y:
0.15x+0.15(-40+x)=7
0.3x=13
x=43 (rounded)
y=-40+43
y=3

Hence, the speed of the boat is 43mph and the speed of the stream is 3mph.
Paul.